OpenStudy (chycora):

5. (1 pt) Baird and Lon earned points for the number of books and magazines they read over the summer. The average of the two boys’ points can be expressed as , where p represents the number of points earned per book. What is the average when p = 3?

OpenStudy (love_ranaa):

19.4 :)

OpenStudy (anonymous):

5.4 * 3 = ?

OpenStudy (anonymous):

@Love_Ranaa Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the user to the answer, so we need his participation. http://openstudy.com/code-of-conduct

OpenStudy (chycora):

16.2

OpenStudy (anonymous):

3.2 * 3 = ?

OpenStudy (chycora):

9.6

OpenStudy (anonymous):

Yes, and what's 16.2 + 5?

OpenStudy (chycora):

21.2

OpenStudy (anonymous):

8 + 9.6 = ?

OpenStudy (chycora):

17.6

OpenStudy (anonymous):

21.2 + 17.6 = ?

OpenStudy (chycora):

38.8

OpenStudy (anonymous):

38.8 / 2 = ?

OpenStudy (anonymous):

That will be your answer.

OpenStudy (chycora):

19.4